OpenStudy (anonymous):

I need to solve this without a calculator 6tan \frac{ 1 }{ 3 }x =3sqrt{3} from \[\left( 0, 2\pi\right)\]

OpenStudy (anonymous):

is it \[\tan(\frac{1}{3}x)\] or \[\tan(\frac{1}{3})x\] ?

OpenStudy (anonymous):

\[6\tan \frac{ 1 }{ 3 }x =3\sqrt{3}\]

OpenStudy (campbell_st):

so you are trying to solve \[6 \tan(\frac{x}{3}) = 3 \sqrt{3}\]

OpenStudy (campbell_st):

divide both sides by 6 what do you get..?

OpenStudy (anonymous):

but is the x inside or outside the tangent? that's kind of important

OpenStudy (campbell_st):

its inside...

OpenStudy (anonymous):

I feel really bad. The tan is supposed to be sin. When i devide by 6 i get \[\sin \frac{ x }{ 3 }=\frac{ \sqrt{3} }{ 2 }\]

OpenStudy (anonymous):

yep, so how do you reverse sin?

OpenStudy (anonymous):

\[\frac{ 1 }{ \sin } or \csc \] ?

OpenStudy (anonymous):

got it thanks sometimes i make these harder than they need to be